# Does provable absence of barren plateaus imply classical simulability? Or, why we need to rethink variational quantum computing

### v2

@@ V.1 Caveats
First and foremost, we would like to highlight the fact that our general arguments are based on intuition gathered from a case-by-case study of widely used circuit architectures and techniques. It was not possible to analyze every work claiming absence of barren plateaus—we studied only those in Table 1—but we highly encourage the community to check if our results are as widely applicable as we believe them to be. Moreover, while we have argued that absence of barren plateaus is linked to the presence of relevant polynomially small subspaces, we do not close the door to the existence of other non-exponential concentration mechanisms. For instance, Claim 1 states for all cases studied that, if $\mathcal{C}\subset\overline{{\rm BP}}$ then $\mathcal{C}\subset\mathcal{C}_{{\rm polySub}}$. We do not claim that $\overline{{\rm BP}}\subset\mathcal{C}_{{\rm polySub}}$. Thus we are not claiming that any non-concentrated loss can always be classically simulated.
@@
In fact, one can construct examples of non-concentrated loss functions that are not classically simulable (see the Appendix for one such example based on cryptographic hardness). Crucially, these examples do not resemble current mainstream variational quantum algorithms but instead draw inspiration from conventional fault-tolerant quantum algorithms for which we expect a superpolynomial quantum advantage to be achievable. In doing so, they break our initial assumption that comparing objects living in exponentially large spaces generically leads to concentrated expectation values, as the circuits used therein are not generic but rather purposely constructed. In fact, textbook quantum algorithms do not suffer from the curse of dimensionality as the exponentially large quantum states are manipulated in a well-thought and orderly manner, rather than a variational one. Nonetheless, as evidenced by our rather contrived examples in the Appendix, it is not obvious how to use these techniques in variational quantum computing as textbook quantum algorithms are only built for very specific tasks.
@@
It is worth noting that our general arguments against conventional variational quantum algorithms require provable absence of barren plateaus, as the proof itself is used in our derivation of the classical simulability techniques. As such, we cannot comment on situations where one can heuristically find large gradients (e.g., via numerics [39, 100, 101]) but where the relevant subspace is not known. Such situations could arise for very special parts of the landscape, such as warm-starts or other smart initialization strategies. Although we believe that a careful analysis of many such cases will highlight that the problem is essentially contained in a polynomially-sized subspace, it could be entirely possible that its determination might be hard, or that even if one finds it, there is no efficient protocol for characterizing the adjoint action of $U(\bm{\theta})$ or for determining the subspace components of $\rho$ and $O$. Alternatively, as suggested by our counterexamples, it might be that the clever initialization explores an exponentially large space in some structured—but unknown—manner [^7] Care must be taken here because, of course, absence of barren plateaus is not a sufficient condition for quantum advantage. Take, for example, ansätze for quantum chemistry with smart initialization techniques. These can be heuristically shown to not exhibit barren plateaus [39], but it has also been noted that the problems where they work are precisely the ones where classical methods perform well and where only a polynomial advantage can be attained [102]. Moreover, it is also well known that initialization techniques can get trapped in local, sub-optimal minima [73, 103, 72, 104]. Hence, more work is needed to understand the wiggle room in these cases. .
@@
Relatedly, let us highlight that, for problems that live in effective but not proper subspaces, classical simulation is possible for $\bm{\theta}\sim\mathcal{P}$ only with high probability. In particular, we can simulate the parts of the loss that are contained within a polynomially-sized subspace and so exhibit large gradients. However, this leaves open the possibility of anomalous parameter values that are useful for training but are not classically simulable. For example, when using a warm start, the initial parameters may not be representative of the whole landscape. Alternatively, one could wander into such regions during training. More concretely, if the subspace is effective (as in Fig. 3(b)), our methods will provide simulation techniques that will only be faithful in the vicinity of the initialization. However, if the relevant sector $\mathcal{B}_{\lambda}$ shifts during training the simulation might cease to be reliable.
@@ V.2 New opportunities
First and foremost, we would like to note that just because some loss function can be classically simulated in polynomial time, this does not mean that it is still practical to do so. In particular, some of the simulation algorithms we have found (see the Appendix) could still be prohibitively expensive to implement. Thus, by embracing the fact that some barren plateau-free loss functions are classically simulable, one could compare the computational complexity of the simulation versus that of estimating the loss function in a quantum computer, and thus potentially find provable polynomial speed-ups such as it is done in Ref. [106].
@@
Third, and as discussed in the Appendix, there appears to be no ‘one simulation algorithm’ to rule them all, as we have found that different loss functions require different simulation algorithms. Hence, given some barren plateau-free model, we envision that asking the question ‘How can we simulate this loss?’ will lead us to new data-driven and quantum-inspired simulation algorithms. In the same line of thought, we think that our quest to better understand what makes an initial state well-aligned with the relevant subspaces, and thus a barren plateau-free loss, can allow us to uncover what are the true resources that make a loss function more difficult to simulate, but also more concentrated. In parallel, as most classical algorithms will have their limits where they particularly excel or perform poorly, it may also turn out that quantum methods where one runs the circuit on the device could still provide the most flexible ‘catch all’ simulation method.
@@
As an example, we refer the reader to Ref. [52], where a connection between computational resources in fermionic linear optic circuits, absence of barren plateaus, and simulability is unraveled. Therein, it was shown that when the quantum resources (measured by fermionic entanglement) are increased, the loss function becomes more concentrated. Simultaneously, even if the state has a large component in the polynomial subspace one still needs to estimate it, and whether such computation can be classically performed will likely determine if the loss is in ${\rm CSIM}$ or ${\rm CSIM}_{\mathrm{QE}}$ (see Claim 2). We thus leave for future work a detailed study of the initial states and measurement operators for which the problems are actually in ${\rm CSIM}$ [^8] Note that this analysis could promote simulations that are available on average, to faithful simulations for all $\bm{\theta}$. For instance, a quantum convolutional neural network with product initial states can be efficiently simulated via tensor networks [88], and hence is actually in ${\rm CSIM}$ $\forall\bm{\theta}$.. We believe that such pursuits will lead to deep insights into architecture-dependent resource theories and the limits of classical simulation algorithms.
@@
Going further, we have noted that problems which are in ${\rm CSIM}_{{\rm QE}}$ but not in ${\rm CSIM}$ are precisely those that are more amenable for implementation in the near term. This is due to the fact that the data acquisition might require shallower circuits than the implementation of $U(\bm{\theta})$, and thus be less impacted by hardware noise. Moreover, optimizing directly on the classical device will generally be much more efficient given the relative speed of classical devices and the possibility of using tools such as automatic differentiation through gradient back-propagation. Indeed, while the loss is defined as a unitary evolution of the state’s projection onto the polynomial subspace, once the data is stored classically, one can potentially analyze it with more general procedures such as sending it through an appropriate classical neural network. This then begs the question of whether the best way to analyze the tomographic data is via a unitary map, or some more general function.
@@
Regarding the utility of variational quantum computing and machine learning for classical data [4, 5, 6, 7, 8], our results should help us better understand if the quantum computer is being used as a meaningful form of enhanced feature-space. Namely, if the problem’s loss function is barren plateau-free, and we can identify how to simulate it, one should ask whether the loss is in ${\rm CSIM}$ or in ${\rm CSIM}_{{\rm QE}}$. Answering these questions could rule out situations where the use of a quantum computer is not necessary. One could further explore the parallels here with recent research into the possibility of classical surrogates for quantum models using classical data [118, 119, 120, 91, 121].
@@ VI Conclusions
Our argument here at times borders on trivial, and may be—in some form, or another—already known to many. However, the connection between the absence of barren plateaus and simulability still has not permeated the field’s Zeitgeist, and it is not uncommon to find claims where the absence of barren plateaus is equated with practical usefulness. In this manner, our conclusions push against current practice in the community and the net result is that variational quantum algorithms need, at best, a serious rethink.
So, where does this leave us? One path forward is to embrace our case-by-case argument as something positive. The fact that a quantum computer is generally not required to implement the parameterized quantum circuit, but might be needed for an initial data acquisition stage, makes such algorithms much easier to implement. To take advantage of this opportunity will require research into how to best perform the data acquisition stage and how to best perform the classical simulation needed to optimize the loss. Moreover, careful analytic work will be required to understand the scaling of the classical simulation algorithms. It may well turn out that simulating the parametrized quantum circuit on quantum hardware enjoys polynomial advantages over our best classical algorithms.
Alternatively, one could rally against our arguments here and strive harder to identify potential avenues for achieving an exponential quantum advantage with parameterized quantum circuits. Of course, there are numerous classically hard quantum circuits and so it’s natural to ask: Can’t this be used to show that there are variational quantum algorithms that are barren plateau-free but not classically simulable? Indeed, we showed that this is possible for a few contrived examples. The key question is whether the non-simulability of quantum circuits can be used more generally to construct useful, trainable, and non-classically simulable variational quantum algorithms. To achieve this we should seek to draw inspiration from conventional fault-tolerant quantum algorithms, and carefully consider how best to manipulate information in the exponentially large spaces in which quantum operators live. Variational quantum computing may still be a viable line of research in this regard but it will require a principled approach and a healthy dose of imagination.
@@
All in all, we can only hope that this perspective article will allow us to pause, reconsider, and continue with a more principled, carefully thought, approach to variational quantum computing.
@@ VII Acknowledgments
We are extremely grateful to Hsin-Yuan Huang for his invaluable contributions to this work. We thank Andrew Sornborger, Lukasz Cincio, Nathan Wiebe, Chae-Yeun Park, Nathan Killoran, Maria Schuld, Xanadu’s Toronto office staff, and the QTML 2023 community for thoughtful and insightful conversations. M.C. acknowledges support from Los Alamos National Laboratory (LANL) ASC Beyond Moore’s Law project. M.L. was supported by the Center for Nonlinear Studies at LANL. M.C., D.G.M., N.L.D. and P.B. were supported by Laboratory Directed Research and Development (LDRD) program of LANL under project numbers 20230527ECR and 20230049DR. Also, N.L.D. acknowledges support from CONICET Argentina, and P.B acknowledges support of DIPC. A.I. acknowledges support by the U.S. Department of Energy (DOE) through a quantum computing program sponsored by the LANL Information Science & Technology Institute and by the U.S. DOE, Office of Science, Office of Advanced Scientific Computing Research, under Computational Partnerships program. E.F. acknowledges the support of the UK department for Business, Energy and Industrial Strategy through the National Quantum Technologies Programme, and the support of an industrial CASE (iCASE) studentship, funded by the Engineering and Physical Sciences Research Council (grant EP/T517665/1), in collaboration with the University of Strathclyde, the National Physical Laboratory, and Quantinuum. E.R.A. acknowledges support from the Walter Burke Institute for Theoretical Physics at Caltech. S.T. and Z.H. acknowledge support from the Sandoz Family Foundation-Monique de Meuron program for Academic Promotion. M.C., Z.H., and E.R.A. thank the organizers of the PennyLane Research Retreat, where part of this work was undertaken, for their hospitality. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (LC). This work was also supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. DOE.
Appendices
@@
Appendix A Barren plateau-free models, polynomial subspaces, and simulation algorithms
@@
We now present an analysis for each architecture and technique in Tables 1 and 2. In all cases we begin by defining the problem class of interest, reviewing the proof of absence of barren plateaus (i.e., determining the measurement operators and initial states leading to non-exponential concentration), and using this information to find the polynomially-large subspace. We then present tomographic techniques that can allow us to obtain the projection of the initial state onto the relevant subspace, and some appropriate simulation algorithms that can be used to classically compute the loss. In particular, here we show that that loss functions in Tables 1 and 2 are in ${\rm CSIM}_{\mathrm{QE}}$.

### v3

@@ V.1 Caveats
First and foremost, we would like to highlight the fact that our general arguments are based on intuition gathered from a case-by-case study of widely used circuit architectures and techniques (although we do refer the reader to [78, 79, 80, 81, 70, 82, 83, 114] for explicit simulation algorithms). It was not possible to analyze every work claiming absence of barren plateaus—we studied only those in Table 1—but we highly encourage the community to check if our results are as widely applicable as we believe them to be. Moreover, while we have argued that absence of barren plateaus is linked to the presence of relevant polynomially small subspaces, we do not close the door to the existence of other non-exponential concentration mechanisms. For instance, Claim 1 states for all cases studied that, if $\mathcal{C}\subset\overline{{\rm BP}}$ then $\mathcal{C}\subset\mathcal{C}_{{\rm polySub}}$. We do not claim that $\overline{{\rm BP}}\subset\mathcal{C}_{{\rm polySub}}$. Thus we are not claiming that any non-concentrated loss can always be classically simulated.
@@
In fact, one can construct examples of non-concentrated loss functions that are not classically simulable (see the Supplemental Information for one such example based on cryptographic hardness). Crucially, these examples do not resemble current mainstream variational quantum algorithms but instead draw inspiration from conventional fault-tolerant quantum algorithms for which we expect a superpolynomial quantum advantage to be achievable. In doing so, they break our initial assumption that comparing objects living in exponentially large spaces generically leads to concentrated expectation values, as the circuits used therein are not generic but rather purposely constructed. In fact, textbook quantum algorithms do not suffer from the curse of dimensionality as the exponentially large quantum states are manipulated in a well-thought and orderly manner, rather than a variational one. Nonetheless, as evidenced by our rather contrived examples in the Supplemental Information, it is not obvious how to use these techniques in variational quantum computing as textbook quantum algorithms are only built for very specific tasks.
Another caveat is that the amount of quantum resources required by problems that fall into $\rm QESIM$} varies on a case-by-case basis. Many cases, as shown in Table 2, require only Pauli measurements and thus the load on the quantum device is very light. Indeed, a universal quantum computer may not be required and rather a more bespoke analogue simulator or quantum experiment may suffice. At the other extreme one could consider a deep quantum circuit with only a single trainable parameter $\phi$ [94]. Such a circuit could be constructed so that it does not have a barren plateau and yet also fall into $\rm QESIM$}, as one can trivially construct a classical surrogate for the landscape by running the circuit in ‘a data collection phase’ at different $\phi$ values. The absence of barren plateau results for small patches of quantum landscapes close to minima [160, 161, 162] which can also be surrogated by running the parametrized quantum circuit and making appropriate measurements [80]. These cases thus strictly align with our claims that problems that can provably avoid barren plateaus fall into $\rm QESIM$} but stretch the original spirit of $\rm QESIM$} by requiring substantially more quantum resources from the quantum device.
@@
It is worth noting that our general arguments against conventional variational quantum algorithms require provable absence of barren plateaus, as the proof itself is used in our derivation of the classical simulability techniques. As such, we cannot comment on situations where one can heuristically find large gradients (e.g., via numerics [41, 163, 164]) but where the relevant subspace is not known. Such situations could arise for very special parts of the landscape, such as warm-starts or other smart initialization strategies. Although we believe that a careful analysis of many such cases will highlight that the problem is essentially contained in a polynomially-sized subspace. For instance, we can refer to the specific case of ADAPT-VQE [41], which has no formal proof of absence of barren plateaus but has been heuristically shown to have large gradients for close to identity initializations. Here, it has been recently shown that Majorana Propagation [84] could efficiently simulate the action of the ADAPT-VQE circuit. Of course, these are preliminary results and there is as of yet no efficient protocol for characterizing the adjoint action of $U(\bm{\theta})$ throughout training and guaranteeing that it will remain in the simulable region. Indeed, as suggested by our counterexamples, it might be that the clever initializations explore an exponentially large space in some structured—but unknown—manner.
@@
Relatedly, let us highlight that, for problems that live in effective but not proper subspaces, classical simulation is possible for $\bm{\theta}\sim\mathcal{P}$ only with high probability. While this can enable for the simulation of randomly chosen points in certain regions of the landscape [80, 162], this need not be necessarily useful for training, as one may be in a non-interesting part of the landscape, or the training might take us towards regions that not classically simulable. More concretely, if the subspace is effective (as in Fig. 3(b)), our methods will provide simulation techniques that will only be faithful in the vicinity of the initialization. However, if the relevant sector $\mathcal{B}_{\lambda}$ shifts during training the simulation might cease to be reliable. The extent to which this is possible is yet to be explored, but recent results have shown that this phenomenon does not seem to occur in quantum convolutional neural networks [78] when benchmarked in standard tests, as there one initializes –and remains during training– in the same classically simulable region of the landscape.
@@ V.2 New opportunities
First and foremost, we would like to note that just because some loss function can be classically simulated in polynomial time, this does not mean that it is still practical to do so. In particular, some of the simulation algorithms we have found (see the Supplemental Information) could still be prohibitively expensive to implement. Thus, by embracing the fact that some barren plateau-free loss functions are classically simulable, one could compare the computational complexity of the simulation versus that of estimating the loss function in a quantum computer, and thus potentially find provable polynomial speed-ups or at least more favorable constant factors with the same polynomial scaling. We refer the reader to Ref. [166, 167, 81] for some examples on this research direction.
@@
Third, and as discussed in the Supplemental Information, there appears to be no ‘one simulation algorithm’ to rule them all, as we have found that different loss functions require different simulation algorithms. Hence, given some barren plateau-free model, we envision that asking the question ‘How can we simulate this loss?’ will lead us to new data-driven and quantum-inspired simulation algorithms. In the same line of thought, we think that our quest to better understand what makes an initial state well-aligned with the relevant subspaces, and thus a barren plateau-free loss, can allow us to uncover what are the true resources that make a loss function more difficult to simulate, but also more concentrated. For instance, the recent work of [179] studied why certain quantum experiments [180] could be efficiently simulated via tensor networks [181]. Therein, the authors showed that the dynamics generated by the quantum circuit ultimately live in a small (identifiable) subspace. In parallel, as most classical algorithms will have their limits where they particularly excel or perform poorly, it may also turn out that quantum methods where one runs the circuit on the device could still provide the most flexible ‘catch all’ simulation method.
@@
As an example, we refer the reader to Ref. [54], where a connection between computational resources in fermionic linear optic circuits, absence of barren plateaus, and simulability is unraveled. Therein, it was shown that when the quantum resources (measured by fermionic entanglement) are increased, the loss function becomes more concentrated. Simultaneously, even if the state has a large component in the polynomial subspace one still needs to estimate it, and whether such computation can be classically performed will likely determine if the loss is in $\rm CSIM$} or $\rm QESIM$} (see Claim 2). We thus leave for future work a detailed study of the initial states and measurement operators for which the problems are actually in $\rm CSIM$}. Note that this analysis could promote simulations that are available on average, to faithful simulations for all $\bm{\theta}$. For instance, a quantum convolutional neural network with product initial states can be efficiently simulated via tensor networks [105], and hence is actually in $\rm CSIM$} $\forall\bm{\theta}$. We believe that such pursuits will lead to deep insights into architecture-dependent resource theories and the limits of classical simulation algorithms.
@@
Going further, we have noted that problems which are in $\rm QESIM$} but not in $\rm CSIM$} could be precisely those that are more amenable for implementation in the near term. This is due to the fact that the data acquisition might require shallower circuits than the implementation of $U(\bm{\theta})$, and thus be less impacted by hardware noise. Moreover, optimizing directly on the classical device will generally be much more efficient given the relative speed of classical devices and the possibility of using tools such as automatic differentiation through gradient back-propagation. Indeed, while the loss is defined as a unitary evolution of the state’s projection onto the polynomial subspace, once the data is stored classically, one can potentially analyze it with more general procedures such as sending it through an appropriate classical neural network. This then begs the question of whether the best way to analyze the tomographic data is via a unitary map, or some more general function.
@@
Regarding the utility of variational quantum computing and machine learning for classical data [4, 5, 6, 7, 8], our results should help us better understand if the quantum computer is being used as a meaningful form of enhanced feature-space. Namely, if the problem’s loss function is barren plateau-free, and we can identify how to simulate it, one should ask whether the loss is in $\rm CSIM$} or in $\rm QESIM$}. Answering these questions could rule out situations where the use of a quantum computer is not necessary. One could further explore the parallels here with recent research into the possibility of classical surrogates for quantum models using classical data [182, 183, 184, 110, 185].
@@ VI Conclusions
Our argument here at times borders on trivial, and may be—in some form, or another—already known to many. However, the connection between the absence of barren plateaus and simulability still has not permeated the field’s Zeitgeist, and it is not uncommon to find claims where the absence of barren plateaus is equated with practical usefulness. In this manner, our conclusions push against current practice in the community and the net result is that variational quantum algorithms need a rethink.
So, where does this leave us? One path forward is to embrace our case-by-case argument as something positive. While some classical simulations have been performed [78, 79, 80, 81, 70, 82, 83, 84], much work remains to be done in this field of “quantum machine learning dequantization”. Moreover, the fact that a quantum computer is generally not required to implement the parameterized quantum circuit, but might be needed for an initial data acquisition stage, makes such algorithms much easier to implement. To take advantage of this opportunity will require research into how to best perform the data acquisition stage and how to best perform the classical simulation needed to optimize the loss. Moreover, careful analytic work will be required to understand the scaling of the classical simulation algorithms. It may well turn out that simulating the parametrized quantum circuit on quantum hardware enjoys polynomial advantages over our best classical algorithms.
Alternatively, one could rally against our arguments here and strive harder to identify potential avenues for achieving a provable exponential quantum advantage with parameterized quantum circuits. Of course, there are numerous classically hard quantum circuits and so it’s natural to ask: Can’t this be used to show that there are variational quantum algorithms that are barren plateau-free but not classically simulable even with data from quantum experiments? Indeed, we showed that this is possible for a few contrived examples. The key question is whether the non-simulability of quantum circuits can be used more generally to construct useful, trainable, and non-classically simulable variational quantum algorithms. To achieve this we should seek to draw inspiration from conventional fault-tolerant quantum algorithms, and carefully consider how best to manipulate information in the exponentially large spaces in which quantum operators live. Variational quantum computing may still be a viable line of research in this regard but it will require a principled approach and a healthy dose of imagination.
@@
We hope this perspective encourages the community to pause, reflect and take a more principled approach to variational quantum computing.
@@ VII Acknowledgments
We are extremely grateful to Hsin-Yuan Huang for his invaluable contributions to this work. We thank Andrew Sornborger, Lukasz Cincio, Nathan Wiebe, Chae-Yeun Park, Nathan Killoran, Maria Schuld, Xanadu’s Toronto office staff, and the QTML 2023 community for thoughtful and insightful conversations. M.C. acknowledges support from Los Alamos National Laboratory (LANL) ASC Beyond Moore’s Law project. M.L. was supported by the Center for Nonlinear Studies at LANL. M.C., D.G.M., N.L.D. and P.B. were supported by Laboratory Directed Research and Development (LDRD) program of LANL under project numbers 20230527ECR and 20230049DR. Also, N.L.D. acknowledges support from CONICET Argentina, and P.B acknowledges support of DIPC. A.I. acknowledges support by the U.S. Department of Energy (DOE) through a quantum computing program sponsored by the LANL Information Science & Technology Institute and by the U.S. DOE, Office of Science, Office of Advanced Scientific Computing Research, under Computational Partnerships program. E.F. acknowledges the support of the UK department for Business, Energy and Industrial Strategy through the National Quantum Technologies Programme, and the support of an industrial CASE (iCASE) studentship, funded by the Engineering and Physical Sciences Research Council (grant EP/T517665/1), in collaboration with the University of Strathclyde, the National Physical Laboratory, and Quantinuum. E.R.A. acknowledges support from the Walter Burke Institute for Theoretical Physics at Caltech. S.T. and Z.H. acknowledge support from the Sandoz Family Foundation-Monique de Meuron program for Academic Promotion. ST further acknowledges the grants for development of new faculty staff, Ratchadaphiseksomphot Fund, Chulalongkorn University [grant number 3230120336 DNS_68_052_2300_012], as well as funding from National Research Council of Thailand (NRCT) [grant number N42A680126]. M.C., Z.H., and E.R.A. thank the organizers of the PennyLane Research Retreat, where part of this work was undertaken, for their hospitality. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center (LC). This work was also supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. DOE.
AUTHOR CONTRIBUTIONS
The project was conceived by MC, ML, and ZH. Theoretical results were derived by MC, NLD, and ERA. All authors contributed to the manuscript review process.
DATA AVAILABILITY
No data was generated.
COMPETING INTERESTS
The authors declare no competing interests.
Supplemental Information for “Does provable absence of barren plateaus imply classical simulability?
@@
Supp. Info. A Barren plateau-free models, polynomial subspaces, and simulation algorithms
@@
We now present an analysis for each architecture and technique in Tables 1 and 2. In all cases we begin by defining the problem class of interest, reviewing the proof of absence of barren plateaus (i.e., determining the measurement operators and initial states leading to non-exponential concentration), and using this information to find the polynomially-large subspace. We then present tomographic techniques that can allow us to obtain the projection of the initial state onto the relevant subspace, and some appropriate simulation algorithms that can be used to classically compute the loss. In particular, here we show that that loss functions in Tables 1 and 2 are in $\rm QESIM$}.